OpenStudy (anonymous):

Match each property with the example. Answer Commutative Read Answer Items for Question 9 Associative Read Answer Items for Question 9 Distributive Read Answer Items for Question 9 Multiplicative of Zero Read Answer Items for Question 9 Multiplicative Identity Read Answer Items for Question 9 Answer A. 5(1) = 5 B. 5(0) = 0 C. 4(x + 5) = 4x + 20 D. (x + 5) + 2 = x + (5 + 2) E. 5 + x = x + 5

jimthompson5910 (jim_thompson5910):

tell me what you got you don't have to answer them all, I just want to see how far you got on this one

OpenStudy (anonymous):

C is distributive property and B is multiplicative of zero... that all?

jimthompson5910 (jim_thompson5910):

C is correct

jimthompson5910 (jim_thompson5910):

B is also correct, nice job so far

jimthompson5910 (jim_thompson5910):

so any ideas on why 5(1) = 5 is true

jimthompson5910 (jim_thompson5910):

which property would you use here

OpenStudy (anonymous):

well i dont know what is Commutative Associative Multiplicative Identity ? is multiplicative identity?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

5(1) = 5 is true by the multiplicative identity

jimthompson5910 (jim_thompson5910):

1*any number = same number

jimthompson5910 (jim_thompson5910):

any number times 1 = same number

jimthompson5910 (jim_thompson5910):

that's what the property is saying

jimthompson5910 (jim_thompson5910):

now how about 5 + x = x + 5

OpenStudy (anonymous):

ohh ok i see... and is it commutative?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

reordering addition/multiplication terms is using the commutative property

jimthompson5910 (jim_thompson5910):

x+y = y+x x * y = y * x

OpenStudy (anonymous):

ok that helped alot thanx.

jimthompson5910 (jim_thompson5910):

yw